A student sets up the following equation to convert a measurement. (The ? stands for a number the student is going to calculate.) Fill in the missing part of this equation ,-10=- 1.4 × 10 - cm
Answer #1

Ans : Here , g/cm^3 has to be converted to g/m^3

1 grams/ cm^3 =======> 1000000 g/m^3

So. -1.4 × 10^4 grams/cm^3 =======> X

Therefore X × 1 = -1.4 × 10^4 × 1000000 grams / m^3

So X = - 1.4 × 10^4 × 10^6 g/ m^3

So , the missing part which must be written in the empty box is 1 × 10^6 to convert g/cm^3 to g/m^3